# A kernel-derived orthogonal basis for spectral functions from Euclidean correlators

### 24 Mar 2026 v1

@@ 1 Introduction
Spectral functions encode essential dynamical information of quantum field theories and many-body systems, including particle excitation spectra, decay rates, and transport coefficients. In lattice QCD and related numerical approaches, such information is accessed indirectly through Euclidean correlation functions. The relation between a Euclidean correlator and the corresponding spectral function is typically expressed as an integral transform with a smooth kernel, whose inversion constitutes a paradigmatic example of an ill-posed problem [1].
A large number of techniques have been developed to address this inverse problem, including Bayesian inference and maximum entropy methods [2, 3, 4, 5, 6], singular value decomposition [7], Padé-based approaches [8], the Backus–Gilbert method [9, 11, 10], and other regularized reconstruction schemes [12, 13, 14, 15, 16, 17]  [^1] For reviews, see, for example, Ref. [1] and references therein.. While these methods have achieved important progress, they often rely on priors, regularization assumptions. Establishing systematic approaches that make minimal assumptions remains a topic of active interest. See for recent progress in this direction, for exmaple, Refs. [18, 19, 20].
@@ 3 Approximating $\tilde{\rho}/\tilde{\omega}$ in an orthogonal basis
Here and in the following, the sum is taken over the repeated indices. Substituting (16) into (17), we obtain
@@ 4 Examples
$\tilde{\rho}_{1}$ is borrowed from [17] with a slight modification to guarantee it is always positive, and inspired from a spectral function related to an "electric field" correlator yielding the momentum-diffusion coefficient of a heavy quark [22, 23, 24]. $\tilde{\rho}_{2}$, inspired in two-dimensional Hubbard model, is taken from Ref. [26], in which the methods based on the maximum entropy, singular value decomposition, sampling, and Padé approximants are compared. $\tilde{\rho}_{3}$ is randomly chosen to have a structure different from $\tilde{\rho}_{1}$ and $\tilde{\rho}_{2}$, and probably it is difficult for many methods to reproduce.
@@
They roughly follow the input spectral function in either cases although one can not guess true input only from them.
@@
which is obtained by substituting (31) into (25), is already very accurate. But it should be noted that each coefficient in (35) is the result of the cancellation among large values of $O(10^{2})$ or $O(10^{3})$ and hence the accuracy in the practical calculation will depend on the errors of $\tilde{C}_{n}^{(m)}$ and the correlation among them. $\Gamma=\lim_{\tilde{\omega}\to 0}{\tilde{\rho}(\tilde{\omega})/\tilde{\omega}}$ is often related to a transport coefficient and is also well reproduced better than 2 % for the expansion using only three basis functions.
@@ 5 Summary
We have illustrated the properties of this expansion using several model spectral functions under idealized conditions. The results demonstrate that global features of spectral functions, as well as low-energy quantities such as transport coefficients, can be reproduced with good accuracy using only a limited number of basis elements, provided the spectral function does not exhibit rapid oscillatory behavior. At the same time, the analysis reveals inherent numerical challenges arising from the exponentially decreasing eigenvalues, which amplify the sensitivity to statistical and systematic uncertainties in practical applications.

### 27 Mar 2026 v2

@@ 1 Introduction
Spectral functions encode essential dynamical information of quantum field theories and many-body systems, including particle excitation spectra, decay rates, and transport coefficients. In lattice QCD and related numerical approaches, such information is accessed indirectly through Euclidean correlation functions. The relation between a Euclidean correlator and the corresponding spectral function is typically expressed as an integral transform with a smooth kernel, whose inversion constitutes a paradigmatic example of an ill-posed problem  [^1] See, for example, [1] and for review [2]..
A large number of techniques have been developed to address this inverse problem, including Bayesian inference and maximum entropy methods [3, 4, 5, 6, 7], singular value decomposition [8], Padé-based approaches [9], the Backus–Gilbert method [10, 12, 11], and other regularized reconstruction schemes [13, 14, 15, 16, 17, 18]. While these methods have achieved important progress, they often rely on priors, regularization assumptions. Establishing systematic approaches that make minimal assumptions remains a topic of active interest. See for recent progress in this and related topics, for example, [19, 20, 21, 22, 23].
@@ 3 Approximating $\tilde{\rho}/\tilde{\omega}$ in an orthogonal basis
Here and thereafter, the sum is taken over repeated indices. Substituting (16) into (17), we obtain
@@ 4 Examples
$\tilde{\rho}_{1}$ is borrowed from [18] with a slight modification to guarantee it is always positive, and inspired from a spectral function related to an "electric field" correlator yielding the momentum-diffusion coefficient of a heavy quark [25, 26, 27]. $\tilde{\rho}_{2}$, inspired in the study of two-dimensional Hubbard model, is taken from Ref. [28], in which the methods based on the maximum entropy, singular value decomposition, sampling, and Padé approximants are compared. $\tilde{\rho}_{3}$ is randomly chosen to have a structure different from $\tilde{\rho}_{1}$ and $\tilde{\rho}_{2}$, and probably it is difficult for many methods to reproduce.
@@
They roughly follow the input spectral function in either cases although one can not guess the true input from them.
@@
which is obtained by substituting (31) into (25), is already very accurate. But it should be noted that each coefficient in (35) is the result of the cancellation among large values of $O(10^{2})$ or $O(10^{3})$ and hence the accuracy in the practical calculation will depend on the errors of $\tilde{C}_{n}^{(m)}$ and the correlation among them. $\Gamma=\lim_{\tilde{\omega}\to 0}{\tilde{\rho}(\tilde{\omega})/\tilde{\omega}}$ is often related to a transport coefficient and is also well reproduced better than 2 % in the expansion using only three basis functions.
@@ 5 Summary
We have illustrated the properties of this expansion using several model spectral functions under idealized conditions. The results demonstrate that global features of spectral functions, as well as low-energy quantities such as transport coefficients, can be reproduced with good accuracy using only a limited number of basis elements even when the spectral function exhibits rapid oscillatory behavior. It is also important to recall that the analysis remains the same even for the spectral function with a finite momentum. At the same time, the analysis reveals inherent numerical challenges arising from the exponentially decreasing eigenvalues, which amplify the sensitivity to statistical and systematic uncertainties in practical applications.